OpenStudy (anonymous):

The stopping distance d for a car after the brake are applied varies directly as the square of the speed r. If a car traveling 70 mph can stop in 270 ft in 270ft, how many feet will it take the same car to stop when it is traveling 40 mph? It will take how many feet?

OpenStudy (amistre64):

d = kr^2

OpenStudy (anonymous):

I put 270 in the twice sry

OpenStudy (amistre64):

270 = k50^2 270/50^2 = k d = 270(40^2)/50^2 maybe :)

OpenStudy (amistre64):

i get 172.8

OpenStudy (anonymous):

i need integer or fraction

OpenStudy (amistre64):

172' 8/10 172' 4/5 172(5)+4 -------- = 864/5 5